Moderately Large Corporation Consolidaed Balance Sheet in thousands except share data) Dec. 31, 2013 Current assets: 1,427 876 s 1,369 Cash and cash equivalents Accounts receivable, net Inventories Prepaid expenses and other current assets Deferred income taxes, net 481 1,489 157 40 4、066 3,137 168 7,371 2,950 2,287 161 5.398 Total current assets Property, plant and equipment, net Other assets TOTAL ASSETS LIABILITIES AND SHAREHOLDERS EQUITY Current liabilities: 242 98 Accounts payable Accrued compensation and related costs Accrued taxes Current portion of long-term debt 429 104 132 89 754 2,630 3,384 563 1,830 Total current liabilities Long-term debt Total liabilities Shareholders equity: Common stock ($0.1 par value)- authorized 4,000,000 shares; issued and outstanding, 3,500,000 Paid-in capital in excess of par 350 2,415 1,222 3,987 350 2,415 241 3,006 Retained earnings Total shareholders equity TOTAL LIABILITIES AND SHAREHOLDERS EQUITY

dos ers ers ers o ces 44 18. 9 6 25 11 32 13 011- $50 57 rze 25-003 112 93 22 201 EST 16 11-005 ser pas- 500 000 | 44 4 247 33 1 250-009 05-004 193 16 125 79-13 ssi 11 ,, dez 11149 258 43 1 3| 865 m 61-4 58 55 50 89-52 22 10-12 29 5 2114 853 1 2 4 ac ed in

1,913 1,687 1338 % 1,691 Operating income Interest expense 1,522 10 1,616 4 64% 11000% Earnings before income taxes 1,512 1,680 588 457% 457% 1,605 Income taxes 529 983 $ 1,092 $ 1,043 4.67% 0.28 Net Earnings Per Common Share Net Earnings Basic Net Earnings Diluted 4.67% 4.67% 0.31 0.25$ 0.26 Weighted average shares outstanding Basic Diluted 3,500 4,000 3,500 4,000 3,500 4,000 0.00% 0.00% 3. Given the income statement for the MLC (Table 4-7) and balance sheet (Table 4-4), answer the following: a. Calculate the following ratios for 2013: operating profit margin, net profit margin, operating return on assets, net return on assets, and return on equity.

Ratio 2013 Operating profit margin- Net profit margin Operating return on assets- Net return on assets- Return on equity b.In a written explanation, deseribe what each of these ratios means and the impact on MLC. Operating profit margin Net profit margin Operating return on assets Net return on assets Return on equity c. In a brief paragraph, describe the overall profitability of the MLC. Identify your concerns.
Answer #1

Part (a)

Tax Rate=(Income tax / Earning Before Income Tax)*100 =(529/1512)*100= 35%

Profit After Tax-983

Interest Net of tax=10(1-35%)=6.5

NOPAT=983+6.5=989.50

1. Operating Profit Ratio= [Net Operating Profit After Tax(NOPAT) / Net Revenue]*100

=(989.50 / 8935)*100

=11.07%

2. Net profit Ratio = [ Net Profit (PAT) / Net Revenue ]*100

=(983/8935)*100

=11.00%

3.Operating Return on Assets = [ Net Operating Profit After Tax(NOPAT) / Average Net Assets]*100

={989.50/[(7371+5398)/2]}*100

=15.50%

4. Net Return on Assets = [ Net Profit After Tax(PAT) / Average Net Assets]*100

={983/[(7371+5398)/2]}*100

=15.40%

5. Return on equity = [ Net Profit After Tax(PAT) / Average Shareholder Funds]*100

={983/[(3384+2392)/2]}*100

=34.04%

Part (b)

1. Operating Profit Ratio=

Net Operating profit of such co in related to revenue is 11.07%

2. Net profit Ratio=

Net profit of such co in related to revenue is 11.0%

3.Operating Return on Assets=

Operationg Profit given by assets used in business is 15.50

4. Net Return on Assets=

Net Profit given by assets used in business is 15.50

5. Return on equity=

Net Profit earned on Shareholders Fund utilised by the company

Part (c)

Growth rate in cost of sales (21.70) tends to be higher than growth rate in Sales (11.56)